# Partial Label Learning with a Partner

### v1

@@ Introduction
where $\hat{\mathbf{b}}=[\hat{b}_{1},\hat{b}_{2},...,\hat{b}_{l}]^{\mathsf{T}}\in\mathbb{R}^{l}$ is the bias term, $\mathbf{1}_{n}\in\mathbb{R}^{n}$ is an all one vectors, $\mathbf{1}_{n\times l}$ is an all one matrix with size $n\times l$ and $\lambda$ is a hyper-parameter trading off these terms. $\|\hat{\mathbf{W}}\|_{F}$ is the Frobenius norm of the weight matrix. $\mathbf{C}\in\mathbb{R}^{n\times l}$ represents non-candidate labeling confidence, which is a temporary variable only used for optimization in the partner classifier. By solving this optimization problem, the partner classifier learns to specify which labels should not be assigned to each sample, improving the overall performance of the mutual supervision process.
The Collaborative Term
@@
The problem in Eq. (5) can be solved via an alternative and iterative manner, which results in an optimal weight matrix $\hat{\mathbf{W}}$ and bias term $\hat{\mathbf{b}}$. For detailed solution process, please refer to the Supplementary. The modeling output $\hat{\mathbf{M}}$ for the training data is
@@
where $\mathcal{T}_{1}$, $\mathcal{T}_{\hat{\mathbf{Y}}}$ are two thresholding operators in element-wise, i.e., $\mathcal{T}_{1}(m):=\min\{1,m\}$ with $m$ being a scalar and $\mathcal{T}_{\hat{\mathbf{Y}}}(m):=\max\{\hat{{y}}_{ij},m\}$. The non-candidate labeling confidence can be further processed as $\hat{\mathbf{Q}}_{1}$ via the blurring mechanism:
$\hat{\mathbf{Q}}_{1}=\phi(e^{k}(1-\hat{\mathbf{P}}_{1}))\odot\mathbf{Y}.$ (8)
For convenience, we transform the non-candidate labeling confidence into labeling confidence in advance, and finally $\hat{\mathbf{Q}}_{1}$ is normalized as $\hat{\mathbf{O}}_{1}$, which is the supervision of the base classifier in the next iteration. For an unseen sample $\mathbf{x}$, suppose the non-candidate labeling confidence vector predicted by the partner classifier in the last iteration is $\hat{\mathbf{p}}^{pt}$. The prediction $y^{\ast}$ of PLCP is
@@
Extensions of PLCP
@@
We can also extend PLCP to a kernel version which extends the feature map to a higher dimensional space or a deep-learning based version which enables deep-learning based methods involved in PLCP. For more details, please refer to the Supplementary.
Experiments
Compared Approaches
@@
To evaluate the effectiveness of PLCP, we couple it with several well-established partial label learning approaches. Suppose $\mathcal{B}$ represents any partial label learning classifier (i.e., the base classifier) and $\mathcal{B}$-PLCP is the $\mathcal{B}$ coupled with PLCP, the performances of $\mathcal{B}$ and $\mathcal{B}$-PLCP are compared to verify the effectiveness of PLCP. In this paper, $\mathcal{B}$ is instantiated by six stand-alone (non-deep) approaches, PL-CL(Jia, Si, and Zhang 2023), PL-AGGD (Wang, Zhang, and Li 2022), SURE (Feng and An 2019), LALO (Feng and An 2018), PL-SVM (Nguyen and Caruana 2008) and PL-KNN (Hüllermeier and Beringer 2005), and two deep-learning based methods PICO (Wang et al. 2022) and PRODEN (Lv et al. 2020). The hyper-parameters of $\mathcal{B}$ are all set according to the original papers.
@@
FG-NET Lost MSRCv2 Mirflickr Soccer Player Yahoo!News
PL-CL 0.072 $\pm$ 0.009 0.710 $\pm$ 0.022 0.469 $\pm$ 0.016 0.647 $\pm$ 0.012 0.534 $\pm$ 0.004 0.618 $\pm$ 0.003
PL-CL-PLCP 0.080 $\pm$ 0.009 $\bullet$ 0.763 $\pm$ 0.020 $\bullet$ 0.493 $\pm$ 0.013 $\bullet$ 0.665 $\pm$ 0.011 $\bullet$ 0.543 $\pm$ 0.002 $\bullet$ 0.625 $\pm$ 0.002 $\bullet$
PL-AGGD 0.063 $\pm$ 0.010 0.690 $\pm$ 0.020 0.451 $\pm$ 0.023 0.610 $\pm$ 0.012 0.521 $\pm$ 0.004 0.605 $\pm$ 0.002
PL-AGGD-PLCP 0.076 $\pm$ 0.010 $\bullet$ 0.717 $\pm$ 0.020 $\bullet$ 0.473 $\pm$ 0.017 $\bullet$ 0.668 $\pm$ 0.014 $\bullet$ 0.534 $\pm$ 0.005 $\bullet$ 0.609 $\pm$ 0.002 $\bullet$
SURE 0.052 $\pm$ 0.007 0.709 $\pm$ 0.022 0.445 $\pm$ 0.022 0.630 $\pm$ 0.022 0.519 $\pm$ 0.004 0.598 $\pm$ 0.002
SURE-PLCP 0.076 $\pm$ 0.011 $\bullet$ 0.719 $\pm$ 0.019 $\bullet$ 0.460 $\pm$ 0.020 $\bullet$ 0.657 $\pm$ 0.020 $\bullet$ 0.527 $\pm$ 0.004 $\bullet$ 0.606 $\pm$ 0.002 $\bullet$
LALO 0.065 $\pm$ 0.010 0.682 $\pm$ 0.019 0.449 $\pm$ 0.016 0.629 $\pm$ 0.016 0.523 $\pm$ 0.003 0.601 $\pm$ 0.003
LALO-PLCP 0.076 $\pm$ 0.010 $\bullet$ 0.701 $\pm$ 0.019 $\bullet$ 0.453 $\pm$ 0.015 $\bullet$ 0.647 $\pm$ 0.018 $\bullet$ 0.529 $\pm$ 0.004 $\bullet$ 0.605 $\pm$ 0.002 $\bullet$
PL-SVM 0.043 $\pm$ 0.008 0.406 $\pm$ 0.033 0.389 $\pm$ 0.029 0.516 $\pm$ 0.022 0.412 $\pm$ 0.006 0.509 $\pm$ 0.006
PL-SVM-PLCP 0.081 $\pm$ 0.011 $\bullet$ 0.688 $\pm$ 0.029 $\bullet$ 0.468 $\pm$ 0.025 $\bullet$ 0.607 $\pm$ 0.023 $\bullet$ 0.526 $\pm$ 0.005 $\bullet$ 0.609 $\pm$ 0.002 $\bullet$
PL-KNN 0.036 $\pm$ 0.006 0.300 $\pm$ 0.018 0.393 $\pm$ 0.014 0.454 $\pm$ 0.016 0.492 $\pm$ 0.003 0.368 $\pm$ 0.004
PL-KNN-PLCP 0.076 $\pm$ 0.009 $\bullet$ 0.662 $\pm$ 0.025 $\bullet$ 0.469 $\pm$ 0.016 $\bullet$ 0.607 $\pm$ 0.023 $\bullet$ 0.523 $\pm$ 0.004 $\bullet$ 0.593 $\pm$ 0.004 $\bullet$
Improvement: PL-CL: 3.61%  PL-AGGD: 5.10 %  SURE: 12.24 %  LALO: 4.01 %  PL-SVM: 39.26 %  PL-KNN: 53.98 %
Table 1: Classification accuracy of each compared approach on the real-world data sets. For any compared approach $\mathcal{B}$, $\bullet$/$\circ$ indicates whether $\mathcal{B}$-PLCP is statistically superior/inferior to $\mathcal{B}$ according to pairwise $t$-test at significance level of 0.05.
@@
Fully Supervised $\mathcal{B}$: 94.91 $\pm$ 0.07 %  $\mathcal{B}$-PLCP: 95.02 $\pm$ 0.03 $\mathcal{B}$: 73.56 $\pm$ 0.10 %  $\mathcal{B}$-PLCP:  90.30 $\pm$ 0.08 %
@@
Fully Supervised $\mathcal{B}$:  90.03 $\pm$ 0.13 $\mathcal{B}$-PLCP: 73.69 $\pm$ 0.14 $\mathcal{B}$: 65.03 $\pm$ 0.35 $\mathcal{B}$-PLCP: 65.52 $\pm$ 0.32 %
Table 2: Classification accuracy of each compared approach on CIFAR-10 and CIFAR-100. For any compared approach $\mathcal{B}$, $\bullet$/$\circ$ indicates whether $\mathcal{B}$-PLCP is statistically superior/inferior to $\mathcal{B}$ according to pairwise $t$-test at significance level of 0.05.
Approaches Data set
FG-NET Lost MSRCv2 Mirflickr Soccer Player Yahoo!News
PL-CL 0.159 $\pm$ 0.016 0.832 $\pm$ 0.019 0.585 $\pm$ 0.012 0.697 $\pm$ 0.019 0.715 $\pm$ 0.001 0.827 $\pm$ 0.003 =
PL-CL-PLCP 0.180 $\pm$ 0.011 $\bullet$ 0.852 $\pm$ 0.011 $\bullet$ 0.638 $\pm$ 0.008 $\bullet$ 0.704 $\pm$ 0.021 $\bullet$ 0.719 $\pm$ 0.002 $\bullet$ 0.829 $\pm$ $0$bullet
PL-AGGD 0.141 $\pm$ 0.012 0.793 $\pm$ 0.020 0.557 $\pm$ 0.015 0.695 $\pm$ 0.015 0.669 $\pm$ 0.003 0.808 $\pm$ 0.005
PL-AGGD-PLCP 0.165 $\pm$ 0.014 $\bullet$ 0.827 $\pm$ 0.019 $\bullet$ 0.640 $\pm$ 0.015 $\bullet$ 0.715 $\pm$ 0.015 $\bullet$ 0.713 $\pm$ 0.003 $\bullet$ 0.831 $\pm$ 0.004 $\bullet$
SURE 0.158 $\pm$ 0.012 0.796 $\pm$ 0.026 0.603 $\pm$ 0.016 0.650 $\pm$ 0.024 0.700 $\pm$ 0.003 0.798 $\pm$ 0.005
SURE-PLCP 0.170 $\pm$ 0.013 $\bullet$ 0.834 $\pm$ 0.024 $\bullet$ 0.621 $\pm$ 0.013 $\bullet$ 0.699 $\pm$ 0.025 $\bullet$ 0.703 $\pm$ 0.003 $\bullet$ 0.827 $\pm$ 0.005 $\bullet$
LALO 0.153 $\pm$ 0.017 0.818 $\pm$ 0.019 0.548 $\pm$ 0.009 0.681 $\pm$ 0.013 0.688 $\pm$ 0.004 0.822 $\pm$ 0.004
LALO-PLCP 0.168 $\pm$ 0.018 $\bullet$ 0.831 $\pm$ 0.019 $\bullet$ 0.620 $\pm$ 0.009 $\bullet$ 0.694 $\pm$ 0.019 $\bullet$ 0.706 $\pm$ 0.004 $\bullet$ 0.827 $\pm$ 0.004 $\bullet$
PL-SVM 0.176 $\pm$ 0.015 0.609 $\pm$ 0.055 0.570 $\pm$ 0.040 0.581 $\pm$ 0.022 0.660 $\pm$ 0.008 0.691 $\pm$ 0.005
PL-SVM-PLCP 0.192 $\pm$ 0.012 $\bullet$ 0.786 $\pm$ 0.032 $\bullet$ 0.639 $\pm$ 0.031 $\bullet$ 0.628 $\pm$ 0.027 $\bullet$ 0.709 $\pm$ 0.006 $\bullet$ 0.821 $\pm$ 0.004 $\bullet$
PL-KNN 0.041 $\pm$ 0.007 0.337 $\pm$ 0.030 0.415 $\pm$ 0.014 0.466 $\pm$ 0.013 0.493 $\pm$ 0.004 0.403 $\pm$ 0.010
PL-KNN-PLCP 0.166 $\pm$ 0.012 $\bullet$ 0.784 $\pm$ 0.031 $\bullet$ 0.635 $\pm$ 0.015 $\bullet$ 0.626 $\pm$ 0.019 $\bullet$ 0.698 $\pm$ 0.004 $\bullet$ 0.790 $\pm$ 0.008 $\bullet$
Table 3: Transductive accuracy of each compared approach on the real-world data sets. For any compared approach $\mathcal{B}$, $\bullet$/$\circ$ indicates whether $\mathcal{B}$-PLCP is statistically superior/inferior to $\mathcal{B}$ according to pairwise $t$-test at significance level of 0.05.

### v3

@@ 2.3 Partner Classifier
where $\hat{\mathbf{b}}=[\hat{b}_{1},\hat{b}_{2},...,\hat{b}_{l}]^{\mathsf{T}}\in\mathbb{R}^{l}$ is the bias term, $\mathbf{1}_{n}\in\mathbb{R}^{n}$ is an all one vectors, $\mathbf{1}_{n\times l}$ is an all one matrix with size $n\times l$ and $\lambda$ is a hyper-parameters trading off these terms. $\|\hat{\mathbf{W}}\|_{F}$ is the Frobenius norm of the weight matrix. $\mathbf{C}\in\mathbb{R}^{n\times l}$ represents non-candidate labeling confidence, which is a temporary variable only used for optimization in the partner classifier. By solving this optimization problem, the partner classifier learns to specify which labels should not be assigned to each sample, improving the overall performance of the mutual supervision process.
2.4 The Collaborative Term
@@
The problem in Eq. (5) can be solved via an alternative and iterative manner in Sec. 2.6, which results in an optimal weight matrix $\hat{\mathbf{W}}$ and bias term $\hat{\mathbf{b}}$. The modeling output $\hat{\mathbf{M}}$ for the training data is
@@
where $\mathcal{T}_{1}$, $\mathcal{T}_{\hat{\mathbf{Y}}}$ are two thresholding operators in element-wise, i.e., $\mathcal{T}_{1}(m):=\min\{1,m\}$ with $m$ being a scalar and $\mathcal{T}_{\hat{\mathbf{Y}}}(m):=\max\{\hat{{y}}_{ij},m\}$. The non-candidate labeling confidence can be further processed as $\hat{\mathbf{O}}_{1}$ via the blurring mechanism:
$\hat{\mathbf{Q}}_{1}=$ $\mathcal{E}(e^{k}(1-\hat{\mathbf{P}}_{1}))\odot\mathbf{Y},$ (8)
$[\hat{\mathbf{O}}_{1}]_{i}$ $=\frac{[\hat{\mathbf{Q}}_{1}]_{i}}{\|[\hat{\mathbf{Q}}_{1}]_{i}\|}.$
For convenience, we transform the non-candidate labeling confidence into labeling confidence in advance, and finally $\hat{\mathbf{Q}}_{1}$ is normalized as $\hat{\mathbf{O}}_{1}$, which is the supervision of the base classifier in the next iteration.
Algorithm 1 The pseudo code of PLCP
1:  Input: A base classifier $\mathcal{B}$ and the partner classifier $\mathcal{P}$, partial label data $\mathcal{D}$, hyper-parameter $\alpha$, $\gamma$ and $\lambda$, temperature parameter $k$, the maximum number of iteration $iter$ and unseen sample $\mathbf{x}^{\ast}$.
2:  Output: The predicted label $y^{\ast}$
3:  Initialize $\mathbf{P}_{0}$, $\hat{\mathbf{P}}_{0}$ and $i=1$.
[…]
5:     Train the base classifier supervised by $\hat{\mathbf{O}}_{i-1}$ and obtain the modeling output $\mathbf{M}_{i}$.
6:     $\mathbf{P}_{i}=\mathcal{T}_{0}\left(\mathcal{T}_{\mathbf{Y}}\left(\alpha\mathbf{P}_{i-1}+(1-\alpha)\mathbf{M}_{i}\right)\right)$.
7:     $\mathbf{Q}_{i}=\mathcal{E}(e^{k}\mathbf{P}_{i})\odot\mathbf{Y}$.
8:     Normalize $\mathbf{Q}_{i}$ and return $\mathbf{O}_{i}$.
9:     Train the partner classifier supervised by $\mathbf{O}_{i}$ and obtain the modeling output $\hat{\mathbf{M}}_{i}$ according to Eq. (6).
10:     $\hat{\mathbf{P}}_{i}=\mathcal{T}_{1}\left(\mathcal{T}_{\mathbf{\hat{Y}}}\left(\alpha\hat{\mathbf{P}}_{i-1}+(1-\alpha)\hat{\mathbf{M}}_{i}\right)\right)$.
11:     $\hat{\mathbf{Q}}_{i}=\mathcal{E}(e^{k}(1-\hat{\mathbf{P}}_{i}))\odot\mathbf{Y}$.
12:     Normalize $\hat{\mathbf{Q}}_{i}$ and return $\hat{\mathbf{O}}_{i}$.
13:     if the stopping criterion is meet then
14:        break;
15:     end if
16:     $i=i+1$.
17:  end while
18:  Return: The predicted label $\mathbf{y}^{\ast}$ according to Eq. (9).
2.5 Stopping Criterion
Through continuous mutual supervision, the base classifier could identify and rectify mislabeled samples. This iterative process of mutual supervision persists for multiple iterations. The iteration ceases when there is a consistent correction of over 5% of mislabeled samples across two consecutive iterations.
For an unseen sample $\mathbf{x}^{\ast}$, suppose the non-candidate labeling confidence vector predicted by the partner classifier in the last iteration is $\hat{\mathbf{p}}^{pt}$. The prediction $y^{\ast}$ of PLCP is
@@ 2.5 Stopping Criterion
Rewrite $\mathbf{P}_{0}=\mathbf{P}$, $\hat{\mathbf{P}}_{0}=\hat{\mathbf{P}}$, $\mathbf{H}_{1}=\mathbf{H}$, $\hat{\mathbf{H}}_{1}=\hat{\mathbf{H}}$ and $\hat{\mathbf{O}}_{0}=\mathbf{P}$, the complete procedure of PLCP is shown in Algorithm 1.
2.6 Numerical Solution of the Partner Classifier
Th optimization problem in Eq. (5) has three variables with different constraints, therefore we can adopt an alternative and iterative optimization to solve it.
2.6.1 Update $\hat{\mathbf{W}}$ and $\hat{\mathbf{b}}$
With $\mathbf{C}$ fixed, the problem w.r.t. $\hat{\mathbf{W}}$ and $\hat{\mathbf{b}}$ can be written as
$\min_{\hat{\mathbf{W}},\hat{\mathbf{b}}}\quad\left\|\mathbf{X}\hat{\mathbf{W}}+\mathbf{1}_{n}\hat{\mathbf{b}}^{\mathsf{T}}-\mathbf{C}\right\|_{F}^{2}+\lambda\left\|\hat{\mathbf{W}}\right\|_{F}^{2},$ (10)
which is a least square problem with the closed-form solution as
$\begin{split}\hat{\mathbf{W}}&=\left(\mathbf{X}^{\mathsf{T}}\mathbf{X}+\lambda\mathbf{I}_{n\times n}\right)\mathbf{X}^{\mathsf{T}}\mathbf{C},\\ \hat{\mathbf{b}}&=\frac{1}{n}\left(\mathbf{C}^{\mathsf{T}}\mathbf{1}_{n}-\hat{\mathbf{W}}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}\mathbf{1}_{n}\right),\end{split}$ (11)
where $\mathbf{I}_{n\times n}$ is the identity matrix with the size $n\times n$.
2.6.2 Update $\mathbf{C}$
With $\hat{\mathbf{W}}$ and $\hat{\mathbf{b}}$ fixed, the $\mathbf{C}$-subproblem can be formulated as
$\begin{split}\min_{\mathbf{C}}&\quad\left\|\mathbf{X}\hat{\mathbf{W}}+\mathbf{1}_{n}\hat{\mathbf{b}}^{\mathsf{T}}-\mathbf{C}\right\|_{F}^{2}+\gamma{\rm tr}\left(\mathbf{O}_{1}\mathbf{C}^{\mathsf{T}}\right)\\ {\rm s.t.}&\quad\hat{\mathbf{Y}}\leq\mathbf{C}\leq\mathbf{1}_{n\times l},\mathbf{C}\mathbf{1}_{l}=(l-1)\mathbf{1}_{n}.\end{split}$ (12)
For simplicity, $\mathbf{O}_{1}$ is written as $\mathbf{O}$ and $\mathbf{J}=\mathbf{X}\hat{\mathbf{W}}+\mathbf{1}_{n}\hat{\mathbf{b}}^{\mathsf{T}}$. Notice that each row of $\mathbf{C}$ is independent to other rows, therefore the problem in Eq. (12) can be solved row by row:
$\begin{split}\min_{\mathbf{C}_{i}}&\quad\mathbf{C}_{i}^{\mathsf{T}}\mathbf{C}_{i}+\left(\gamma\mathbf{O}_{i}-2\mathbf{J}_{i}\right)^{\mathsf{T}}\mathbf{C}_{i}\\ {\rm s.t.}&\quad\hat{\mathbf{Y}}_{i}\leq\mathbf{C}_{i}\leq\mathbf{1}_{n},\mathbf{C}_{i}\mathbf{1}_{l}=l-1.\end{split}$ (13)
The problem in Eq. (13) is a standard Quadratic Programming (QP) problem, which can be solved by off-the-shelf QP tools.
2.7 Extensions of PLCP
@@
We can also extend PLCP to a kernel version which extends the feature map to a higher dimensional space or a deep-learning based version which enables deep-learning based methods involved in PLCP.
2.7.1 Kernel Extension
The linear classifier in Eq. (5) may lack some ability to tackle the complex relationships among the data. Therefore, we extend it to a kernel version. Denote $\phi(\cdot):\mathbb{R}^{q}\rightarrow\mathbb{R}^{h}$ the feature mapping that maps the feature space to some higher dimension space with $h$ dimensions, then we can rewrite Eq. (10) as
$\begin{split}\min_{\hat{\mathbf{W}},\hat{\mathbf{b}}}\quad&\|\mathbf{Z}\|_{F}^{2}+\lambda\|\hat{\mathbf{W}}\|_{F}^{2}\\ {\rm s.t.}\quad&\mathbf{Z}=\mathbf{\Phi}\hat{\mathbf{W}}+\mathbf{1}_{n}\hat{\mathbf{b}}^{\mathsf{T}}-\mathbf{C},\end{split}$ (14)
where $\mathbf{\Phi}=[\phi(\mathbf{x}_{1}),\phi(\mathbf{x}_{2}),...,\phi(\mathbf{x}_{n})]$. The Lagrangian function of this problem is
$\begin{split}&\mathcal{L}(\hat{\mathbf{W}},\hat{\mathbf{b}},\mathbf{Z},\mathbf{A})=\|\mathbf{Z}\|_{F}^{2}+\lambda\|\hat{\mathbf{W}}\|_{F}^{2}\\ &-{\rm tr}(\mathbf{A}^{\mathsf{T}}(\mathbf{\Phi}\hat{\mathbf{W}}+\mathbf{1}_{n}\hat{\mathbf{b}}^{\mathsf{T}}-\mathbf{C}-\mathbf{Z})),\end{split}$ (15)
where $\mathbf{A}=[a_{ij}]_{n\times l}\in\mathbb{R}^{n\times l}$ stores the Lagrange multipliers. Following the KKT conditions:
$\begin{split}&\frac{\partial\mathcal{L}}{\partial\hat{\mathbf{W}}}=2\lambda\hat{\mathbf{W}}-\mathbf{\Phi}^{\mathsf{T}}\mathbf{A}=0,\frac{\partial\mathcal{L}}{\partial\hat{\mathbf{b}}}=\mathbf{A}^{\mathsf{T}}\mathbf{1}_{n}=0,\\ &\frac{\partial\mathcal{L}}{\partial\mathbf{Z}}=2\mathbf{Z}+\mathbf{A}=0,\frac{\partial\mathcal{L}}{\partial\mathbf{A}}=\mathbf{\Phi}\hat{\mathbf{W}}+\mathbf{1}_{n}\hat{\mathbf{b}}^{\mathsf{T}}-\mathbf{C}-\mathbf{Z}=0,\end{split}$ (16)
we have the corresponding solution:
$\begin{split}&\hat{\mathbf{W}}=\frac{\mathbf{\Phi}^{\mathsf{T}}\mathbf{A}}{2\lambda},\mathbf{A}=\left(\frac{1}{2\lambda}\mathbf{K}+\frac{1}{2}\mathbf{I}_{n\times n}\right)^{-1}\left(\mathbf{C}-\mathbf{1}_{n}\hat{\mathbf{b}}^{\mathsf{T}}\right),\\ &\mathbf{s}=\mathbf{1}_{n}^{\mathsf{T}}\left(\frac{1}{2\lambda}\mathbf{K}+\frac{1}{2}\mathbf{I}_{n\times n}\right)^{-1},\hat{\mathbf{b}}^{\mathsf{T}}=\frac{\mathbf{s}\mathbf{C}}{\mathbf{s1}_{n}},\end{split}$ (17)
where $\mathbf{K}=\mathbf{\Phi}\mathbf{\Phi}^{\mathsf{T}}$ is the kernel matrix with its element $k_{ij}=\phi(\mathbf{x}_{i})^{\mathsf{T}}\phi(\mathbf{x}_{j})=\mathcal{K}(\mathbf{x}_{i},\mathbf{x}_{j})$ based on the kernel function $\mathcal{K}(\cdot,\cdot)$. For PLCP, Gaussian function $\mathcal{K}(\mathbf{x}_{i},\mathbf{x}_{j})={\rm exp}(-\frac{\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}}{2\sigma^{2}}$) is adopted as the kernel function with $\sigma$ to the average distance of all pairs of training samples. The modeling output is denoted by
$\hat{\mathbf{M}}=\frac{1}{2\lambda}\mathbf{KA}+\mathbf{1}_{n}\hat{\mathbf{b}}^{\mathsf{T}}.$ (18)
2.7.2 Deep Learning Extension
PLCP can be also extended to a deep-learning version to further demonstrate its universality and effectiveness. The model of the majority of the deep-learning based methods usually contains a feature encoder followed by a prediction head, which predicts the labeling confidence of each sample. Denote a model in $\mathcal{B}$ with such architecture $g(\cdot)$, specifically, an additional model $\hat{g}(\cdot)$ with the same architecture as $g(\cdot)$ is introduced as the partner classifier, which predicts the non-candidate labeling confidence of each sample. Given an example $\mathbf{x}$ and its candidate label vector $\mathbf{y}$, the non-candidate loss is defined as
$\mathcal{L}_{com}=-\sum_{i=1}^{l}(1-y_{i})\log(\hat{g}_{i}(\mathbf{x})).$ (19)
Afterwards, a collaborative loss is designed to link the two models:
$\mathcal{L}_{col}=\sum_{i=1}^{l}v_{i}\hat{v}_{i}.$ (20)
Here, $\mathbf{v}=[v_{1},v_{2},...,v_{l}]$ and $\hat{\mathbf{v}}=[\hat{v}_{1},\hat{v}_{2},...,\hat{v}_{l}]$ are two blurred predictions of $g(\mathbf{x})$ and $\hat{g}(\mathbf{x})$ respectively, where
$\begin{split}\mathbf{v}&=\frac{\mathcal{E}(e^{k}g(\mathbf{x}))\odot\mathbf{y}}{\|\mathcal{E}(e^{k}g(\mathbf{x}))\odot\mathbf{y}\|}\\ \hat{\mathbf{v}}&=\mathbf{1}_{l}^{\mathsf{T}}-\frac{\mathcal{E}(e^{k}(1-\hat{g}(\mathbf{x})))\odot\mathbf{y}}{\|\mathcal{E}(e^{k}(1-\hat{g}(\mathbf{x})))\odot\mathbf{y}\|}.\end{split}$ (21)
The overall loss function is:
$\mathcal{L}=\mathcal{L}_{ori}+\mathcal{L}_{com}+\mu\mathcal{L}_{col},$ (22)
where $\mathcal{L}_{ori}$ is the original loss function of $\mathcal{B}$ and $\mu$ is a hyper-parameter.
3 Experiments
3.1 Compared Approaches
@@
To evaluate the effectiveness of PLCP, we couple it with several well-established partial label learning approaches. Suppose $\mathcal{B}$ represents any partial label learning classifier (i.e., the base classifier) and $\mathcal{B}$-PLCP is the $\mathcal{B}$ coupled with PLCP, the performances of $\mathcal{B}$ and $\mathcal{B}$-PLCP are compared to verify the effectiveness of PLCP. In this paper, $\mathcal{B}$ is instantiated by six stand-alone (non-deep) approaches, PL-CL[17], PL-AGGD [16], SURE [20], LALO [15], PL-SVM [22] and PL-KNN [21], and two deep-learning based methods PICO [23] and PRODEN [24]. The hyper-parameters of $\mathcal{B}$ are all set according to the original papers.
TABLE I: Characteristics of the real-world data sets.
Data set #examples #features #labels average #labels
FG-NET [14] 1002 262 78 7.48
Lost [18] 1122 108 16 2.23
MSRCv2 [25] 1758 48 23 3.16
Mirflickr [26] 2780 1536 14 2.76
Soccer Player [11] 17472 279 171 2.09
Yahoo!News [12] 22991 163 219 1.91
TABLE II: Classification accuracy of each compared approach on the real-world data sets. For any compared approach $\mathcal{B}$, $\bullet$/$\circ$ indicates whether $\mathcal{B}$-PLCP is statistically superior/inferior to $\mathcal{B}$ according to pairwise $t$-test at significance level of 0.05.
@@
FG-NET Lost MSRCv2 Mirflickr Soccer Player Yahoo!News FG-NET(MAE3) FG-NET(MAE5)
PL-CL 0.072 $\pm$ 0.009 0.710 $\pm$ 0.022 0.469 $\pm$ 0.016 0.647 $\pm$ 0.012 0.534 $\pm$ 0.004 0.618 $\pm$ 0.003 0.433 $\pm$ 0.022 0.575 $\pm$ 0.015
PL-CL-PLCP 0.081 $\pm$ 0.006 $\bullet$ 0.762 $\pm$ 0.017 $\bullet$ 0.496 $\pm$ 0.010 $\bullet$ 0.667 $\pm$ 0.010 $\bullet$ 0.545 $\pm$ 0.001 $\bullet$ 0.628 $\pm$ 0.002 $\bullet$ 0.450 $\pm$ 0.013 $\bullet$ 0.595 $\pm$ 0.006 $\bullet$
PL-AGGD 0.063 $\pm$ 0.010 0.690 $\pm$ 0.020 0.451 $\pm$ 0.023 0.610 $\pm$ 0.012 0.521 $\pm$ 0.004 0.605 $\pm$ 0.002 0.418 $\pm$ 0.020 0.562 $\pm$ 0.020
PL-AGGD-PLCP 0.076 $\pm$ 0.009 $\bullet$ 0.717 $\pm$ 0.020 $\bullet$ 0.474 $\pm$ 0.015 $\bullet$ 0.668 $\pm$ 0.014 $\bullet$ 0.534 $\pm$ 0.003 $\bullet$ 0.611 $\pm$ 0.002 $\bullet$ 0.442 $\pm$ 0.018 $\bullet$ 0.582 $\pm$ 0.011 $\bullet$
SURE 0.052 $\pm$ 0.007 0.709 $\pm$ 0.022 0.445 $\pm$ 0.022 0.630 $\pm$ 0.022 0.519 $\pm$ 0.004 0.598 $\pm$ 0.002 0.356 $\pm$ 0.020 0.494 $\pm$ 0.021
SURE-PLCP 0.077 $\pm$ 0.009 $\bullet$ 0.719 $\pm$ 0.021 $\bullet$ 0.460 $\pm$ 0.018 $\bullet$ 0.657 $\pm$ 0.020 $\bullet$ 0.526 $\pm$ 0.003 $\bullet$ 0.608 $\pm$ 0.002 $\bullet$ 0.443 $\pm$ 0.015 $\bullet$ 0.584 $\pm$ 0.012 $\bullet$
LALO 0.065 $\pm$ 0.010 0.682 $\pm$ 0.019 0.449 $\pm$ 0.016 0.629 $\pm$ 0.016 0.523 $\pm$ 0.003 0.601 $\pm$ 0.003 0.422 $\pm$ 0.023 0.566 $\pm$ 0.015
LALO-PLCP 0.076 $\pm$ 0.010 $\bullet$ 0.703 $\pm$ 0.017 $\bullet$ 0.456 $\pm$ 0.018 $\bullet$ 0.647 $\pm$ 0.018 $\bullet$ 0.529 $\pm$ 0.004 $\bullet$ 0.605 $\pm$ 0.002 $\bullet$ 0.444 $\pm$ 0.015 $\bullet$ 0.585 $\pm$ 0.015 $\bullet$
PL-SVM 0.043 $\pm$ 0.008 0.406 $\pm$ 0.033 0.389 $\pm$ 0.029 0.516 $\pm$ 0.022 0.412 $\pm$ 0.006 0.509 $\pm$ 0.006 0.314 $\pm$ 0.019 0.445 $\pm$ 0.016
PL-SVM-PLCP 0.081 $\pm$ 0.010 $\bullet$ 0.692 $\pm$ 0.024 $\bullet$ 0.468 $\pm$ 0.021 $\bullet$ 0.611 $\pm$ 0.018 $\bullet$ 0.526 $\pm$ 0.006 $\bullet$ 0.609 $\pm$ 0.002 $\bullet$ 0.439 $\pm$ 0.017 $\bullet$ 0.584 $\pm$ 0.015 $\bullet$
PL-KNN 0.036 $\pm$ 0.006 0.300 $\pm$ 0.018 0.393 $\pm$ 0.014 0.454 $\pm$ 0.016 0.492 $\pm$ 0.003 0.368 $\pm$ 0.004 0.288 $\pm$ 0.014 0.440 $\pm$ 0.017
PL-KNN-PLCP 0.077 $\pm$ 0.006 $\bullet$ 0.668 $\pm$ 0.019 $\bullet$ 0.469 $\pm$ 0.016 $\bullet$ 0.608 $\pm$ 0.026 $\bullet$ 0.523 $\pm$ 0.003 $\bullet$ 0.597 $\pm$ 0.002 $\bullet$ 0.434 $\pm$ 0.020 $\bullet$ 0.578 $\pm$ 0.014 $\bullet$
Average Improvement Ratio: PL-CL: 3.61%  PL-AGGD: 5.10 %  SURE: 12.24 %  LALO: 4.01 %  PL-SVM: 39.26 %  PL-KNN: 53.98 %
TABLE III: Classification accuracy of each compared approach on CIFAR-10 and CIFAR-100. For any compared approach $\mathcal{B}$, $\bullet$/$\circ$ indicates whether $\mathcal{B}$-PLCP is statistically superior/inferior to $\mathcal{B}$ according to pairwise $t$-test at significance level of 0.05.
@@
Fully Supervised $\mathcal{B}$: 94.91 $\pm$ 0.07 %  $\mathcal{B}$-PLCP: 95.02 $\pm$ 0.03 $\mathcal{B}$: 73.56 $\pm$ 0.10 %  $\mathcal{B}$-PLCP:  73.69 $\pm$ 0.14 %
@@
Fully Supervised $\mathcal{B}$:  90.03 $\pm$ 0.13 $\mathcal{B}$-PLCP: 90.30 $\pm$ 0.08 $\mathcal{B}$: 65.03 $\pm$ 0.35 $\mathcal{B}$-PLCP: 65.52 $\pm$ 0.32 %
TABLE IV: Ablation study of PLCP coupled with PL-AGGD on the real-world data sets. $\bullet$/$\circ$ indicates whether PL-AGGD-PLCP is statistically superior/inferior to its degenerated version according to pairwise $t$-test at significance level of 0.05.